Suppose it is known that the population of the community in problem 1 is 10000 after 3 years.


1
Expert's answer
2021-11-08T06:22:59-0500

Let "P_0" Β be the population of a community

The population of a community is known to increase at a rate proportional to the number of people present at a time "t"


"\dfrac{dP}{dt}=kP, k=constant"

Then


"\dfrac{dP}{P}=kdt"

Integrate both sides


"\int\dfrac{dP}{P}=\int kdt"

"\ln(|P)|=kt+\ln C"

"P(t)=Ce^{kt}"


Using the initial condition


"P(t)=P_0e^{kt}"


Given that an initial population "P_0" has doubled in 5 years


"P(5)=P_0e^{k(5)}=2P_0"

"e^{5k}=2"

"k=0.2\ln 2"

"P(t)=P_0\cdot2^{0.2t}"

Given "P(3)=10000."

i)


"10000=P_02^{0.2(3)}"

"P_0=\dfrac{10000}{2^{0.6}}"

"P_0=6598\ people"

ii)


"P(10)=P_0\cdot2^{0.2(10)}"

"P(10)=4P_0"

"P(10)=10000(2)^{3.4}"

"P(10)=26390\ people"

iii)


"\dfrac{dP}{dt}|_{t=10}=0.2\ln 2\cdot P(10)"

"\dfrac{dP}{dt}|_{t=10}=0.2\ln 2\cdot (10000(2)^{3.4})"

"\dfrac{dP}{dt}|_{t=10}=3658.45\ people/year"